What will be the missing digit in 6,10,24,?,185Plz help me out

Grade:12th pass

1 Answers

Malay Damani
5 years ago
6 * 1 + 4 = 10
10 * 1.5 + 9 = 24
24 * 2 + 16 = 64
64 * 2.5 + 25 = 185
185 * 3 + 36 = 591
therfore the general solution of equation is a+n/2+(n+1)2
therefore missing digit is 64